Jolly Travel Agency specializes in flights between Toronto and Jamaica. It books passengers on Burlington Air. Jollys fixedCase 4: Jollys variable costs are $35 per ticket. It receives $55 commission per ticket from Burlington Air. It charges itsCalculate the number of tickets Jolly must sell each month to (a) break even and (b) make a target operating income of $18,00
Answer #1

Break even number of units = total fixed costs/contribution margin per unit

Target operating income =(desired profit + fixed costs)/contribution margin per unit

Contribution margin per unit = revenue per unit- variable costs per unit

1

Break even units =29500/(1600*10%-42)

= 250 units

Target income =(18000+29500)/118

= 402.54 units

2.break even = 29500/(160-35) = 236

To meet operating income =380 tickets

3. Break even = 29500/20 = 1475 tickets

Operating income = 1045 tickets

Lower commission increased break even point

4.break even = 29500/25 = 1180 tickets

Operating income = 1900 tickets

IncreaseD CM

Lowers